# TSPO: Temporal Sampling Policy Optimization for Long-form Video Language Understanding

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant progress in vision-language tasks, yet they still face challenges when processing long-duration video inputs. The limitation arises from MLLMs’ context limit and training costs, necessitating sparse frame sampling before feeding videos into MLLMs. Existing video MLLMs adopt training-free uniform sampling or keyframe search, which may miss critical events or be constrained by the pre-trained models’ event understanding capabilities. Meanwhile, building a training-based method remains challenging due to the unsupervised and non-differentiable nature of sparse frame sampling. To address these problems, we propose Temporal Sampling Policy Optimization (TSPO), advancing MLLMs’ long-form video-language understanding via reinforcement learning. Specifically, we first propose a trainable event-aware temporal agent, which captures event-query correlation for performing probabilistic keyframe selection. Then, we propose the TSPO reinforcement learning paradigm, which models keyframe selection and language generation as a joint decision-making process, enabling end-to-end group relative optimization with efficient rule-based rewards. Furthermore, for the TSPO’s training, we propose a long video training data construction pipeline with comprehensive temporal data and video Needle-in-a-Haystack data. Finally, we incorporate rule-based answering accuracy and temporal locating reward mechanisms to optimize the temporal sampling policy. Comprehensive experiments show that our TSPO achieves state-of-the-art performance across multiple long video understanding benchmarks, and shows transferable ability across different cutting-edge Video-MLLMs. Our code is available at https://github.com/Hui-design/TSPO
@@ Introduction
Multimodal Large Language Models (MLLMs) have achieved significant progress in various vision-language tasks, such as image captioning, visual question answering, OCR, etc. They typically extract visual information as visual tokens into the Large Language Models (LLMs) for open-world understanding. As a natural extension, video-based MLLMs (Video-MLLMs)  (Zhang et al. 2024d; Shen et al. 2024; Bai et al. 2025) have attracted great attention, where videos contain more complex temporal and visual information, bringing more significant challenges.
Figure 1: Illustrations of the video understanding task and different frame sampling methods: (a) uniform sampling suffers from irrelevant frame selection, (b) training-free keyframe extraction fails to capture the essential “assist” semantics, while (c) our method achieves enhanced event localization and response accuracy by exploring and optimizing the keyframe sampling policy for Video-MLLMs.
@@
Existing MLLMs are compelled to employ sparse frame sampling when dealing with videos  (Zhang et al. 2024d; Shen et al. 2024; Kim et al. 2024; Xu et al. 2024; Liu et al. 2024c). Notably, different sampling strategies may yield substantially divergent video interpretations, as highlighted in prior works (Hu et al. 2025a). The core challenge mainly lies in determining the optimal frame sampling strategy that maximizes MLLMs’ video comprehension accuracy while minimizing computational overhead. Most existing Video-MLLM approaches, such as LLaVA-Video (Zhang et al. 2024d) and Qwen2.5vl (Bai et al. 2025), simply perform uniform frame sampling, which results in a considerable loss of valuable information that is relevant to queries. Recently, some studies (Hu et al. 2025a; Shen et al. 2024; Tang et al. 2025) focus on exploring training-free keyframe extraction approaches. For instance, LongVU (Shen et al. 2024) identifies cross-frame distinct frames by leveraging pre-trained feature extractors such as DINOv2-1B (Oquab et al. 2023). CoS (Hu et al. 2025a) employs LLaVA-1.5-13B (Liu et al. 2024a) to filter query-relevant frames for inputting into Video-MLLM, which incurs significant computational costs. Without training optimization, training-free methods are limited by the cross-modal event understanding capabilities of pre-trained keyframe selectors and may incur more significant computational overhead during inference. These limitations lead to a question: Can we develop a trainable sparse frame sampling approach for reliable and efficient long-video language understanding?
@@
However, there exist two fundamental challenges to obtaining a trainable temporal sampling approach for Video-MLLMs: (1) Unsupervised nature: frame-level annotations are generally unavailable in general video understanding training (Zhang et al. 2024d; Bai et al. 2025), resulting in a lack of precise localization guidance. (2) Non-differentiability: frame sampling is a discrete subset selection problem, where the output consists of frame indices rather than continuous variables, making it difficult to optimize via backpropagation in Supervised Fine-Tuning (SFT).
@@
Based on the above analyses and inspired by the progress of Deepseek-R1 (DeepSeek-AI et al. 2025; Shao et al. 2024) in enhancing MLLM reasoning through Group Relative Policy Optimization (GRPO), we propose Temporal Sampling Policy Optimization (TSPO) to learn an optimal keyframe selection strategy for Video-MLLMs. It novelly models keyframe selection and language generation as a joint decision-making process, performing end-to-end GRPO optimization of the temporal agent through rule-based rewards. Specifically, a trainable temporal sampler is first modeled as a decision agent to capture event-query correlation for keyframe probability estimation, which also maintains structural simplicity instead of using other heavy MLLMs like (Hu et al. 2025a, b). Furthermore, for TSPO’s training, we propose a long video training data construction pipeline with comprehensive temporal data for general video understanding and video Needle-in-a-Haystack data for long-range temporal localization. In the reinforcement learning-based temporal sampling policy optimization, we establish efficient rule-based answering accuracy and coarse-level temporal locating reward mechanisms that optimize the temporal agent to maximize the expected reward by choosing critical frames for queries adaptively.
Comprehensive experiments show that TSPO achieves state-of-the-art performance across multiple long video understanding benchmarks. Our contributions are as follows:
@@
  • We propose the Temporal Sampling Policy Optimization algorithm, which models keyframe selection and language generation as a joint decision-making process, performing end-to-end group relative policy optimization through rule-based rewards. This effectively tackles the unsupervised and non-differentiable challenge of sparse frame sampling in Video-MLLMs.
@@
  • Extensive experiments on multiple long video understanding benchmarks with different types of MLLMs as baseline models demonstrate the effectiveness and strong generalization of our TSPO method.
@@
Multimodal Large Language Models (MLLMs) have demonstrated significant progress in vision-language tasks, yet they still face challenges when processing long-duration videos with extremely long context. Previous works (Xu et al. 2024; Liu et al. 2024b; Lin et al. 2023; Cheng et al. 2024; Zhang et al. 2024d, b) often first employ uniform frame sampling and then perform token compression to reduce the length of the context. LLaVA-Video (Zhang et al. 2024d) and SlowFast-LLaVA (Xu et al. 2024) utilize spatial and temporal pooling techniques to decrease the number of tokens. Beyond uniform sampling, recent works are exploring keyframe search methods (Ye et al. 2025; Guo et al. 2025; Wang et al. 2023; Shen et al. 2024; Hu et al. 2025a). For instance, LongVU (Shen et al. 2024) identifies cross-frame distinct frames by leveraging robust feature extractors such as DINOv2 (Oquab et al. 2023) and further discards tokens that exhibit minimal feature differences between the current frame and those of the previous frames. More recently, Chain-of-Shot (Hu et al. 2025a) leverages off-the-shelf multimodal large models, such as LLaVA-1.5 (Liu et al. 2024a), to select task-relevant shots and task-irrelevant shots. However, due to the inherently non-differentiable and unsupervised nature of keyframe sampling in general video understanding tasks, these methods opt for a training-free approach, which cannot be further optimized. In this paper, we propose a learnable temporal agent and present the TSPO algorithm to optimize keyframe sampling.
@@
Reinforcement Learning (RL) is often used in post-training of large language models to align with human preferences, and more recently, to enhance multi-modal capabilities (Wang et al. 2024b; Zhang et al. 2024c; DeepSeek-AI et al. 2025). RLHF(Ouyang et al. 2022) utilizes human feedback to train a reward model, treating the LLM as a policy network and optimizing it using PPO (Schulman et al. 2017). DPO (Rafailov et al. 2024) simplifies RLHF by directly constructing preference data without requiring a reward model. LLaVA-Hound-DPO (Zhang et al. 2024c) and TPO (Li et al. 2025) construct temporal preference data and optimize the LLM within the DPO framework. More recently, Deepseek-R1 (DeepSeek-AI et al. 2025) has garnered significant attention by using the GRPO (Shao et al. 2024) algorithm to achieve robust reasoning capabilities. Video-R1 (Feng et al. 2025) novelly extends GRPO to the more complex video reasoning domain, and proposes large-scale video reasoning datasets. However, these reinforcement learning approaches focus solely on optimizing the reasoning abilities of LLMs. In contrast, our TSPO takes a novel perspective by modeling the discrete keyframe selection and language generation as a unified decision-making process, directly addressing the most challenging issue in current long video understanding: the extremely long context problem.
@@
As shown in Fig. 2, the temporal agent takes CLIP (Radford et al. 2021) frame-level visual features $\mathbf{F}_{f}\in\mathbb{R}^{T_{c}\times D}$ and text features $\mathbf{F}_{t}\in\mathbb{R}^{1\times D}$ as the inputs, where $T_{c}$ denotes the candidate frame number and D denotes the feature dimension. The visual features are then enhanced with event perception capabilities through local window attention (Pu et al. 2024). In detail, the visual features are projected to query, key, and value embeddings through linear layers. Then the attention is restricted to a local window of length $w$ centered at the current frame, augmented with sinusoidal positional embeddings (Vaswani et al. 2017) to learn intra-event dependencies and temporal awareness. This leads to the refined event representation $\mathbf{F}_{e}\in\mathbb{R}^{T_{c}\times D}$, and the cosine similarity $\mathrm{Sim}_{event}(\mathbf{F}_{e},\mathbf{F}_{t})$ is then computed to capture event-text alignment. To strengthen temporal localization robustness, frame-level similarity $\mathrm{Sim}_{frame}(\mathbf{F}_{f},\mathbf{F}_{t})$ between visual features $F_{v}$ and text features $F_{t}$ is concurrently calculated. The final cross-modal similarity $S\in\mathbb{R}^{T_{c}}$ is derived through the fusion of the two scores:
@@
In the reinforcement learning framework, the agent state is defined by the input long video V and the text instruction Q. The agent action corresponds to keyframe selection, outputting the selected indices $\mathcal{I}=\{i_{1},i_{2},..,i_{T_{s}}\}$ and the corresponding probability $\mathcal{P}=\{p_{1},p_{2},..,p_{T_{s}}\}$. To enable diverse action generation for RL exploration (Wei et al. 2023; Cui et al. 2023), our method employs the Gumbel-Softmax (Jang, Gu, and Poole 2016) operator:
@@
In addition, the selected indices $\mathcal{I}$ are used to obtain the sparse frames, which serve as inputs to the Video-MLLMs.
@@
Vanilla Group Relative Policy Optimization (GRPO). Deepseek-R1 (DeepSeek-AI et al. 2025) proposes the GRPO (Shao et al. 2024) algorithm, which foregoes the critic model and estimates the baseline from rule-based rewards instead. Specifically, for each question q from the training set Q, a group of language outputs is sampled $\{\textbf{o}_{1},\textbf{o}_{2},\cdots,\textbf{o}_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model $\pi_{\theta}$ by maximizing:
@@
where $\mathbb{D}_{KL}$ is the unbiased estimator (Schulman 2020) for KL divergence and $\beta$ is a hyper-parameter to balance the weights. $\pi_{ref}$ is the reference model, typically a LLM that has undergone large-scale Supervised Fine-Tuning (SFT). $A_{i}$ is the relative advantage, computed using a group of rewards $\{r_{1},r_{2},\ldots,r_{G}\}$ corresponding to the group outputs (DeepSeek-AI et al. 2025).
@@
$\begin{split}\mathcal{J}_{tspo}(\theta)&=\mathbb{E}_{\textbf{q}\sim\textbf{Q},\textbf{v}\sim\textbf{V},\{\textbf{o}_{i},\textbf{V}_{s}\}\sim\pi_{old}(\textbf{O}\,|\,\mathbf{q},\mathbf{V}_{c})}\\ \frac{1}{G}\sum_{i=1}^{G}&\frac{\pi_{l}(\textbf{o}_{i}\,|\,\mathbf{q},\mathbf{V}_{s},\mathbf{V}_{c})\cdot\pi_{ts}(\textbf{V}_{s}\,|\,\mathbf{q},\mathbf{V}_{c})}{\pi_{{l}_{old}}(\textbf{o}_{i}\,|\,\mathbf{q},\mathbf{V}_{s},\mathbf{V}_{c})\cdot\pi_{{ts}_{old}}(\textbf{V}_{s}\,|\,\mathbf{q},\mathbf{V}_{c})}A_{i}\\ &-\beta\cdot\mathbb{D}_{KL}\left(\pi_{\theta}||\pi_{ref}\right),\end{split}$ (7)
Considering the extremely long context problem is more significant for the current long video understanding model, we maintain focus on optimizing the Temporal Sampler while preserving the strong prior of language generation capabilities. We employ a pre-trained Video-MLLM and keep it frozen, thereby ensuring that:
@@
Therefore, our TSPO objective can be simplified to optimize only the temporal agent, formulated as follows:
@@
GPT-4o (Hurst et al. 2024) - - Uniform 66.7 64.6 65.3 71.9 -
GPT-4V (OpenAI 2023) - - Uniform 61.3 49.2 53.5 59.9 -
Gemini-1.5-Flash (Team et al. 2023) - - Uniform 61.6 - 61.1 70.3 -
Gemini-1.5-Pro (Team et al. 2023) - - Uniform 64 - 67.4 75.0 33.1
Video-LLaVA (Lin et al. 2023) 8 7B Uniform - 36.2 - 39.9 -
Oryx-1.5 (Liu et al. 2025) 64 7B Uniform 56.3 - 51.2 58.8 -
LLaVA-Onevision (Li et al. 2024) 32 7B Uniform 56.4 64.7 46.7 58.2 -
NVILA (Liu et al. 2024d) 1024 7B Uniform 57.7 70.1 54.8 64.2 -
Apollo (Zohar et al. 2024) 2FPS 7B Uniform 58.5 68.7 - 61.3 -
mPLUG-Owl3 (Ye et al. 2024) 128 7B Uniform 59.7 70.0 50.1 59.3 43.5
LongVU (Shen et al. 2024) 1FPS 7B DINOv2-1B - 65.4 - 60.6 -
MLLM-VFS (Hu et al. 2025b) 32 7B MLLM-1.5B 57.0 - 51.9 58.7 -
LLaVA-Video-7B (Zhang et al. 2024d) 64 7B Uniform 58.2 70.8 - 63.3 -
LLaVA-Video-7B+TPO (Li et al. 2025) 64 7B Uniform 60.1 71.1 55.4 65.6 -
LLaVA-Video-7B+CoS (Hu et al. 2025a) 64 7B MLLM-13B 58.9 71.4 53.8 64.4 -
LLaVA-Video-7B+AKS (Tang et al. 2025) 64 7B BLIP-0.5B 62.7 - 54.0 65.3 -
LLaVA-Video-7B* (Zhang et al. 2024d) 64 7B Uniform 58.9 70.3 53.6 64.4 40.2
@@
Table 1: Comparison results on four widely recognized long video understanding benchmarks, where our method achieves state-of-the-art performances with significant accuracy gain. “*” denotes our reproduced results under 64 frames. The first three benchmarks are evaluated using lmms-eval (Zhang et al. 2024a), and LVBench is tested using its own evaluation protocol.
@@
  • LongVideoBench (Wu et al. 2024). We evaluate the validation set with 1,337 videos (avg. 12min), following standard academic protocols (Zhang et al. 2024d).
  • MLVU (Zhou et al. 2024). The video length ranges from 3 minutes to 2 hours. We evaluate on the ”M-Avg” portion of the ”Dev” split, following (Zhang et al. 2024d).
@@
  • LVBench (Wang et al. 2024c) is an extremely long video benchmark, with an average video length of 4,101 seconds—4 times longer than VideoMME.
Implementation Details. The model was trained on 8 NVIDIA A800 80GB GPUs with a single epoch, using a learning rate of $5\times 10^{-4}$ and a batch size of 1. The temporal agent is built upon a frozen CLIP-Large model (400M parameters) and incorporates  only 3.5M learnable parameters. For the Video-MLLM that guides TSPO training, we adopt LLaVA-Video (Zhang et al. 2024d) with its parameters kept frozen. The number of candidate frames ($T_{c}$) is set to 1 FPS, while the selected frame count ($T_{s}$) is set to 64 during inference and 16 during training. The window size $w$ is set to 12, $\tau$ is set at 0.025, and $\alpha$ is set to 0.4. To ensure reproducibility during inference, deterministic predictions were enforced by removing the Gumbel noise.
@@
Selector Parameters. Experimental comparisons show that our lightweight selector (Temporal Agent) achieves both parameter efficiency and superior model performance. Compared to LongVU’s (Shen et al. 2024) 1B-parameter DINOV2 (Oquab et al. 2023) selector, our method achieves a 4.9% absolute accuracy improvement on VideoMME and 10.9% on MLVU. Against Chain-of-Shot (Hu et al. 2025a)’s 13B-scale MLLM selector, our solution outperforms by 6.0% on LongVideoBench and 4.9% on MLVU.
@@
Transferring TSPO to Other Video-MLLMs. Although our method is developed based on LLaVA-Video as the Video-MLLM backbone, we explore an efficient “one-model for all” paradigm (Liu et al. 2023; Cheng et al. 2025) by transferring our learned temporal agent from LLaVA-Video-7B to other Video-MLLMs without extra training, including Qwen2VL (Wang et al. 2024a) / Qwen2.5VL-7B (Bai et al. 2025), and also extending it to LLaVA-Video-72B. As shown in Tab. 2, our method demonstrates notable generalization capability: on LongVideoBench, it achieves an average  4.5% improvement; On the MLVU dataset, the average improvement is 6.3%, with Qwen2.5VL achieving a notably higher gain of 9.2%. This cross-architecture performance verifies the generalizability of our approach.

### v4

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant progress in vision-language tasks, yet they still face challenges when processing long-duration video inputs. The limitation arises from MLLMs’ context limit and training costs, necessitating sparse frame sampling before feeding videos into MLLMs. However, building a trainable sampling method remains challenging due to the unsupervised and non-differentiable nature of sparse frame sampling in Video-MLLMs. To address these problems, we propose Temporal Sampling Policy Optimization (TSPO), advancing MLLMs’ long-form video-language understanding via reinforcement learning. Specifically, we first propose a trainable event-aware temporal agent, which captures event-query correlation for performing probabilistic keyframe selection. Then, we propose the TSPO reinforcement learning paradigm, which models keyframe selection and language generation as a joint decision-making process, enabling end-to-end group relative optimization for the temporal sampling policy. Furthermore, we propose a dual-style long video training data construction pipeline, balancing comprehensive temporal understanding and key segment localization. Finally, we incorporate rule-based answering accuracy and temporal locating reward mechanisms to optimize the temporal sampling policy. Comprehensive experiments show that our TSPO achieves state-of-the-art performance across multiple long video understanding benchmarks, and shows transferable ability across different cutting-edge Video-MLLMs.
Codehttps://github.com/Hui-design/TSPO
@@ Introduction
Multimodal Large Language Models (MLLMs) have achieved significant progress in various vision-language tasks, such as image captioning, visual question answering, OCR, etc. They typically extract visual information as visual tokens into the Large Language Models (LLMs) for open-world understanding. As a natural extension, video-based MLLMs (Video-MLLMs)  (llavavideo; shen2024longvu; qwen2.5vl) have attracted great attention, where videos contain more complex temporal and visual information, bringing more significant challenges.
Figure 1: Illustrations of different frame sampling methods: Training-free uniform sampling (a) and keyframe search (b) select unsatisfactory frames, while our method (c) explores and optimizes the temporal sampling policy that leads to the correct answer in an end-to-end training manner.
@@
Existing MLLMs are compelled to employ sparse frame sampling when dealing with videos  (llavavideo; shen2024longvu; kim2024salova; xu2024slowfast; liu2024ppllava). The core challenge mainly lies in determining the optimal frame sampling strategy that maximizes MLLMs’ video comprehension accuracy while minimizing computational overhead. Most existing Video-MLLM approaches, such as LLaVA-Video (llavavideo) and Qwen2.5VL (qwen2.5vl), simply perform uniform frame sampling, which often misses key information that is relevant to queries, as shown in Fig. 1. Recently, some studies (hu2025cos; shen2024longvu; AKS) focus on exploring training-free keyframe extraction approaches. For instance, LongVU (shen2024longvu) identifies cross-frame distinct frames by leveraging pre-trained feature extractors such as DINOv2-1B (oquab2023dinov2). CoS (hu2025cos) employs LLaVA-1.5-13B (llava) to filter query-relevant frames for inputting into Video-MLLM, which incurs significant computational costs. Without training optimization, training-free methods are limited by the cross-modal event understanding capabilities of pre-trained keyframe selectors and may incur more computation during inference. These limitations lead to a question: Can we develop a trainable sparse frame sampling approach for reliable and efficient long-video language understanding?
@@
However, there exist two fundamental challenges to obtaining a trainable temporal sampling approach for Video-MLLMs: (1) Unsupervised nature: frame-level annotations are generally unavailable in general video understanding training (llavavideo; qwen2.5vl), resulting in a lack of precise localization guidance. (2) Non-differentiability: frame sampling is a discrete subset selection problem, where the output consists of frame indices rather than continuous variables, making it difficult to optimize via backpropagation in Supervised Fine-Tuning (SFT).
@@
Based on the above analyses and inspired by the progress of Deepseek-R1 (deepseekr1; deepseekmath) in enhancing MLLM reasoning through Group Relative Policy Optimization (GRPO), we propose Temporal Sampling Policy Optimization (TSPO) to explore and optimize keyframe selection strategy for Video-MLLMs. It novelly models keyframe selection and language generation as a joint decision-making process, performing end-to-end GRPO optimization of the temporal agent through rule-based rewards. Specifically, a trainable temporal sampler is first modeled as a decision agent to capture event-query correlation for keyframe probability estimation, which also maintains structural simplicity instead of using other heavy MLLMs like (hu2025cos; MLLM-VFS). Furthermore, for TSPO’s training, we propose a long video training data construction pipeline with comprehensive temporal data for general video understanding and video Needle-in-a-Haystack data for long-range temporal localization. In the reinforcement learning-based temporal sampling policy optimization, we establish efficient rule-based answering accuracy and coarse-level temporal locating reward mechanisms that optimize the temporal agent to maximize the expected reward by choosing critical frames for queries adaptively.
Extensive experiments demonstrate the effectiveness and strong generalization of our TSPO method, achieving average performance gains of 4.3% on LLaVA-Video and 6.1% on Qwen2.5-VL. Our contributions are as follows:
@@
  • We propose the Temporal Sampling Policy Optimization algorithm, which models keyframe selection and language generation as a joint decision-making process, performing end-to-end group relative optimization for the temporal sampling policy. This effectively tackles the unsupervised and non-differentiable challenge of sparse frame sampling in Video-MLLMs.
@@
  • Our TSPO achieves state-of-the-art performance across multiple general long video understanding benchmarks, and shows strong transferable ability across different cutting-edge Video-MLLMs.
@@
Multimodal Large Language Models (MLLMs) have demonstrated significant progress in vision-language tasks, yet they still face challenges when processing long-duration videos with extremely long context. Previous works (xu2024procedure; xu2025uni_fineparser; xu2025TransferableUAL; xu2025BFSTAL; xu2024slowfast; stllm; videllava; cheng2024videollama2; llavavideo; longva) often employ uniform frame sampling or perform token compression to reduce the length of the context. LLaVA-Video (llavavideo) and SlowFast-LLaVA (xu2024slowfast) utilize spatial and temporal pooling techniques to decrease the number of tokens. Beyond uniform sampling, recent works are exploring keyframe search methods (T-star; guo2025logic; wang2023vaquita; shen2024longvu; hu2025cos). For instance, LongVU (shen2024longvu) identifies cross-frame distinct frames by leveraging robust feature extractors such as DINOv2 (oquab2023dinov2) and further discards tokens that exhibit minimal feature differences between the current frame and those of the previous frames. More recently, Chain-of-Shot (hu2025cos) leverages off-the-shelf multimodal large models, such as LLaVA-1.5 (llava), to select task-relevant shots and task-irrelevant shots. However, due to the inherently non-differentiable and unsupervised nature of keyframe sampling in general video understanding tasks, these methods opt for a training-free approach, which cannot be further optimized. In this paper, we propose a learnable temporal agent and present the TSPO algorithm to optimize keyframe sampling.
@@
Reinforcement Learning (RL) is often used in post-training of LLMs to align with human preferences. To enhance multi-modal capabilities (MPO; llavahoundDPO; deepseekr1). RLHF(RLHF) utilizes human feedback to train a reward model, treating the LLM as a policy network and optimizing it using PPO (ppo). DPO (DPO) simplifies RLHF by directly constructing preference data without requiring a reward model. LLaVA-Hound-DPO (llavahoundDPO) and TPO (TPO) construct temporal preference data and optimize the LLM within the DPO framework. More recently, Deepseek-R1 (deepseekr1) has garnered significant attention by using the GRPO (deepseekmath) algorithm to achieve robust reasoning capabilities. However, these reinforcement learning approaches focus solely on optimizing the reasoning abilities of LLMs. In contrast, our TSPO takes a novel perspective by modeling the discrete keyframe selection and language generation as a unified decision-making process, directly addressing the most challenging issue in current long video understanding: the extremely long context problem.
@@
As shown in Fig. 2, the temporal agent takes CLIP (clip) frame-level visual features $\mathbf{F}_{f}\in\mathbb{R}^{T_{c}\times D}$ and text features $\mathbf{F}_{t}\in\mathbb{R}^{1\times D}$ as the inputs, where $T_{c}$ denotes the candidate frame number and D denotes the feature dimension. The visual features are then enhanced with event perception capabilities through local window attention (prompt-enhanced). Then the attention is restricted to a local window of length $w$ centered at the current frame, augmented with sinusoidal positional embeddings (attention), and projected by an MLP to learn intra-event dependencies and temporal awareness. This leads to the refined event representation $\mathbf{F}_{e}\in\mathbb{R}^{T_{c}\times D}$, and the cosine similarity $\mathrm{Sim}_{event}(\mathbf{F}_{e},\mathbf{F}_{t})$ is then computed to capture event-text alignment. To strengthen temporal localization robustness, frame-level similarity $\mathrm{Sim}_{frame}(\mathbf{F}_{f},\mathbf{F}_{t})$ between visual features $F_{v}$ and text features $F_{t}$ is concurrently calculated. The final cross-modal similarity $S\in\mathbb{R}^{T_{c}}$ is derived through the fusion of the two scores:
@@
In the reinforcement learning framework, the agent state is defined by the input long video V and the text instruction Q. The agent action corresponds to keyframe selection, outputting the selected indices $\mathcal{I}=\{i_{1},i_{2},..,i_{T_{s}}\}$ and the corresponding probability $\mathcal{P}=\{p_{1},p_{2},..,p_{T_{s}}\}$. To enable diverse action generation for RL exploration (GRANSAC; RDD), our method employs the Gumbel-Softmax (gumbel) operator:
@@
Temperature annealing. A high temperature is used early in training to encourage exploration, and it is gradually reduced to a low temperature to converge to key segments. Then, the selected frames indices $\mathcal{I}$ are used to obtain the sparse frames, which serve as inputs to the Video-MLLMs.
@@
Vanilla Group Relative Policy Optimization (GRPO). Deepseek-R1 (deepseekr1) proposes the GRPO (deepseekmath) algorithm, which foregoes the critic model and estimates the baseline from rule-based rewards instead. Specifically, for each question q from the training set Q, a group of language outputs is sampled $\{\textbf{o}_{1},\textbf{o}_{2},\cdots,\textbf{o}_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model $\pi_{\theta}$ by maximizing:
@@
where $\mathbb{D}_{KL}$ is the unbiased estimator (kl_approx) for KL divergence and $\beta$ is a hyper-parameter to balance the weights. $\pi_{ref}$ is the reference model, typically a LLM that has undergone large-scale Supervised Fine-Tuning (SFT). $A_{i}$ is the relative advantage, computed using a group of rewards $\{r_{1},r_{2},\ldots,r_{G}\}$ corresponding to the group outputs (deepseekr1).
@@
$\begin{split}\mathcal{J}_{tspo}(\theta)&=\mathbb{E}_{\textbf{q}\sim\textbf{Q},\textbf{v}\sim\textbf{V},\{\textbf{o}_{i},\textbf{V}_{s}\}\sim\pi_{old}(\textbf{O}\,|\,\mathbf{q},\mathbf{V}_{c})}\\ \frac{1}{G}\sum_{i=1}^{G}&\frac{\pi_{l}(\textbf{o}_{i}\,|\,\mathbf{q},\mathbf{V}_{s},\mathbf{V}_{c})\cdot\pi_{ts}(\textbf{V}_{s}\,|\,\mathbf{q},\mathbf{V}_{c})}{\pi_{{l}_{old}}(\textbf{o}_{i}\,|\,\mathbf{q},\mathbf{V}_{s},\mathbf{V}_{c})\cdot\pi_{{ts}_{old}}(\textbf{V}_{s}\,|\,\mathbf{q},\mathbf{V}_{c})}A_{i}\\ &-\beta\cdot\mathbb{D}_{KL}\left(\pi_{\theta}||\pi_{ref}\right).\end{split}$ (7)
Considering the extremely long context problem is more significant for the current long video understanding model, we maintain focus on optimizing the Temporal Sampler while preserving the strong prior of language generation capabilities. We employ a pre-trained MLLM and keep it frozen, thereby ensuring that:
@@
Notably, the MLLM has been SFT-trained on LLaVA-Video-178K (our source dataset) with uniform 32 frames and thus can answer questions well when correct keyframes are selected. Therefore, our TSPO objective can be simplified to optimize only the temporal agent as follows:
@@
GPT-4o (gpt4o) - - Uniform 66.7 64.6 65.3 71.9 -
GPT-4V (gpt4v) - - Uniform 61.3 49.2 53.5 59.9 -
Gemini-1.5-Flash (gemini) - - Uniform 61.6 - 61.1 70.3 -
Gemini-1.5-Pro (gemini) - - Uniform 64 - 67.4 75.0 33.1
Video-LLaVA (videllava) 8 7B Uniform - 36.2 - 39.9 -
Oryx-1.5 (orxy1.5) 64 7B Uniform 56.3 - 51.2 58.8 -
LLaVA-Onevision (Llava-onevision) 32 7B Uniform 56.4 64.7 46.7 58.2 -
NVILA (liu2024nvila) 1024 7B Uniform 57.7 70.1 54.8 64.2 -
Apollo (apollo) 2FPS 7B Uniform 58.5 68.7 - 61.3 -
mPLUG-Owl3 (mplug3) 128 7B Uniform 59.7 70.0 50.1 59.3 43.5
LongVU (shen2024longvu) 1FPS 7B DINOv2-1B - 65.4 - 60.6 -
MLLM-VFS (MLLM-VFS) 32 7B MLLM-1.5B 57.0 - 51.9 58.7 -
LLaVA-Video-7B (llavavideo) 64 7B Uniform 58.2 70.8 - 63.3 -
LLaVA-Video-7B+TPO (TPO) 64 7B Uniform 60.1 71.1 55.4 65.6 -
LLaVA-Video-7B+CoS (hu2025cos) 64 7B MLLM-13B 58.9 71.4 53.8 64.4 -
LLaVA-Video-7B+AKS (AKS) 64 7B BLIP-0.5B 62.7 - 54.0 65.3 -
LLaVA-Video-7B* (llavavideo) 64 7B Uniform 58.9 70.3 53.6 64.4 40.2
@@
Table 1: Comparison results on four widely recognized long video understanding benchmarks, where our method achieves state-of-the-art performances with significant accuracy gain. “*” denotes our reproduced results under 64 frames. The first three benchmarks are evaluated using lmms-eval (lmmeval), and LVBench is tested using its own evaluation protocol.
@@
  • LongVideoBench (wu2024longvideobench). We evaluate the validation set with 1,337 videos (avg. 12min), following standard academic protocols (llavavideo).
  • MLVU (zhou2024mlvu). The video length ranges from 3 minutes to 2 hours. We evaluate on the ”M-Avg” portion of the ”Dev” split, following (llavavideo).
@@
  • LVBench (wang2024lvbench) is an extremely long video benchmark, with an average video length of 4,101 seconds—4 times longer than VideoMME.
Implementation Details. The model was trained on 8 NVIDIA A800 80GB GPUs with a single epoch, using a learning rate of $5\times 10^{-4}$ and a batch size of 1. The temporal agent is built upon a frozen CLIP-Large model (400M parameters) and incorporates  only 3.5M learnable parameters. For the Video-MLLM that guides TSPO training, we adopt LLaVA-Video (llavavideo) with its parameters kept frozen. The number of candidate frames ($T_{c}$) is set to 1 FPS, while the selected frame count ($T_{s}$) is set to 64 during inference and 16 during training. The window size $w$ is set to 12, $\tau$ is set at 0.025 and annealed to 0.01. To ensure reproducibility during inference, deterministic predictions were enforced by removing the Gumbel noise.
@@
Selector Parameters. Experimental comparisons show that our lightweight selector (Temporal Agent) achieves both parameter efficiency and superior model performance. Compared to LongVU’s (shen2024longvu) 1B-parameter DINOV2 (oquab2023dinov2) selector, our method achieves a 4.9% absolute accuracy improvement on VideoMME and 10.9% on MLVU. Against Chain-of-Shot (hu2025cos)’s 13B-scale MLLM selector, our solution outperforms by 6.0% on LongVideoBench and 4.9% on MLVU.
@@
Transferring TSPO to Other Video-MLLMs. Although our method is developed based on LLaVA-Video as the Video-MLLM backbone, we explore an efficient “one-model for all” paradigm (oneforall; onemodelforall) by transferring our learned temporal agent from LLaVA-Video-7B to other Video-MLLMs without extra training, including Qwen2VL (wang2024qwen2vl) / Qwen2.5VL-7B (qwen2.5vl), and also extending it to LLaVA-Video-72B. As shown in Tab. 2, our method demonstrates notable generalization capability: on LongVideoBench, it achieves an average  4.5% improvement; On the MLVU dataset, the average improvement is 6.3%, with Qwen2.5VL achieving a notably higher gain of 9.2%. This cross-architecture performance verifies the generalizability of our approach.